2) Emerson mapped three genes in maize: one for green ( V −) or yellow seedling ( vv ); one for rough leaf Gl −) or glossy leaf gl gl ); and one for normal fertility ( Va −) or variable fertility va va He made trihybrids of genotype V Gl Va v gl va and crossed these to v gl va v gl va He analyzed the progeny to map the genes. Answer the following questions analyzing the three point test cross data provided below. 1. What is the gene order on the chromosome? 2. What are the recombination frequencies of the gene pairs? 3. Is the frequency of the double crossovers consistent with independence of the single crossovers?Emersons Three-Point Test-Cross Analysis Table 5.3 vgl va/vgl va yellow, glossy, variable Parental cross: V GI Va/N GIVa Gre
Answer #1

Answer:

1).

The order of genes = V—Gl---Va

Gene map: V------18.32 m.u.----------Gl----------13.64 m.u.-----------Va

2).

Recombination frequency between V & Gl = 18.32%

Recombination frequency between V & Va = 28.93%

Recombination frequency between Gl & Va= 13.64%

Explanation:

Hint: Non-recombinant (parental) combinations are produced more in number than recombinants. The parental (non-recombinant) genotypes is V Gl Va / v gl va

1).

If single crossover occurs between V & Gl

Normal combination: V Gl / v gl

After crossover: V gl / v Gl

V gl progeny=7+62=69

v Gl progeny = 4+60=64

Total this progeny = 133

The recombination frequency between V & Gl = (number of recombinants/Total progeny) 100

RF = (133/726)100 = 18.32%

2).

If single crossover occurs between Gl & Va

Normal combination: Gl Va / gl va

After crossover: Gl va / gl Va

Gl va progeny= 40+4=44

gl Va progeny = 48+7=55

Total this progeny = 99

The recombination frequency between Gl & Va = (number of recombinants/Total progeny) 100

RF = (99/726)100 = 13.64%

3).

If single crossover occurs between V & Va

Normal combination: V Va / v va

After crossover: V va / v Va

V va progeny= 40+62=102

v Va progeny = 48+60=108

Total this progeny = 210

The recombination frequency between V & Va = (number of recombinants/Total progeny) 100

RF = (210/726)100 = 28.93%

Recombination frequency (%) = Distance between the genes (mu)

v--------13.19mu------ct----6.42mu----cv

The order of genes = V—Gl---Va

Gene map: V------18.32 m.u.----------Gl----------13.64 m.u.-----------Va